OpenStudy (anonymous):

In a quadrilateral ABCD, the measure of angle A is 2x, the measure of angle B is 3x-15(degrees), the measure of C is 4x-90(degrees), and the measure of angle D is x+15(degrees). Find the value of X, Is ABCD a rectangle? What is the average of the four angles?

OpenStudy (anonymous):

what do you know about the sum of internal angles of a quadrilateral?

OpenStudy (anonymous):

x=45(degrees) no abcd is not a rectangle average of angles is=90

OpenStudy (anonymous):

not much the whole thing equals 360 degrees

OpenStudy (anonymous):

right, so the sum equals 360. angles A,B,C,D are given. add up all the angles and equate them to 360 post your equation here.

OpenStudy (anonymous):

hint: A, B, C, D being angles, A+B+C+D = 360

OpenStudy (anonymous):

(2x+3x-15+4x-90+x+15=360) = (10x-90=360) x=45.

OpenStudy (anonymous):

right. x = 45. What is the requirement of a quadrilateral to be a rectangle?

OpenStudy (anonymous):

ok, so I was doing something like 2x=360 and then trying to solve from there. totally wrong. all quadrilaterals can be solved like this?

OpenStudy (anonymous):

all the angle measurements have to be 90 degrees right

OpenStudy (anonymous):

yes. so what is the measure of angle A, angle B , angle C and angle D?

OpenStudy (anonymous):

a= 90 b=120 c=90 d=60 average=90 =]

OpenStudy (anonymous):

why do you care about the average?

OpenStudy (anonymous):

ah, okay its asked in the problem. average of the angles of a quadrilateral is always 90, no matter what the angles are.

OpenStudy (anonymous):

because a quadrilateral has 4 internal angles and they add up to 360 degrees. so you always end up with 360/4 as the average angle.

OpenStudy (anonymous):

okay so what is your conclusion? Is this a rectangle or not?

OpenStudy (anonymous):

not a rectangle

OpenStudy (anonymous):

WHAT IS THE DISCUSSION TOPIC?

OpenStudy (anonymous):

very good.

OpenStudy (anonymous):

cool. thanks everybody!

OpenStudy (anonymous):

np

OpenStudy (yuki):

woo